OpenStudy (btsmith101):

Add and simplify 8√x^2y+3√x^2y+√16x^2y

OpenStudy (mathstudent55):

Is this the correct problem? \(8\sqrt{x^2y} + 3\sqrt{x^2y} + \sqrt{16x^2y}\)

OpenStudy (lapinblanc):

If you take the square root of the 16, because it is under the root sign, all of the values under the root signs will be exactly the same. This means you can add together all of the constants in front of them and you answer will be \[(all.constants.added.together)\sqrt{x^{2}y}\]